Drilling operations with Texas Brine continue. Another stage of casing has been installed and the concrete has been poured. It must set for a 24 hr. period.
OEP continues to urge residents, those that have evacuated and those that remain home, to check their homes for any changes such as cracks, ground holes, sinking, etc. Anything unusual or out of the ordinary should be reported to OEP immediately. All reports will be forwarded to DNR for proper investigation.
OEP continues to monitor Tropical Storm Isaac. Preliminary preparations are being made in the event that the tract continues to move in a westerly direction.